# A Generative Framework for Predictive Modeling of Multiple Chronic Conditions Using Graph Variational Autoencoder and Bandit-Optimized Graph Neural Network

### v1

@@ 1 Introduction
Recent advances in the application of machine learning to healthcare have demonstrated significant potential in disease progression prediction and treatment optimization. Carvajal Rico et al. [29] have further advanced the application of GNNs in healthcare by incorporating Laplacian regularization. This technique enhances the learning process by ensuring similar nodes have similar representations, thereby improving the prediction accuracy of models dealing with MCC. Baucum et al. [30] introduced transitional variational autoencoders (tVAEs) to improve reinforcement learning (RL) agents’ performance by generating realistic patient condition trajectories, specifically enhancing the training of medication dosage policies through on-policy RL in intensive care settings. Kmetzsch et al. [31] applied Supervised Variational Autoencoders to predict disease progression scores using multimodal imaging and microRNA data, demonstrating improved accuracy and interpretability in patient outcome predictions. Moreover, Sun et al. [32] highlighted the challenges of using machine learning for disease prediction due to the reliance on abundant manually labeled EMR data, which is often insufficient for rare diseases. Their innovative model leverages external knowledge bases to augment the EMR data and uses GNNs to create highly representative node embeddings for patients, diseases, and symptoms. By aggregating information from connected neighbor nodes, their neural graph encoder can generate embeddings that capture knowledge from both data sources, allowing accurate prediction of general and rare diseases.
@@
In this work, we propose an innovative framework combining a GVAE, a Laplacian regularized graph neural network (LR-GNN), and contextual bandit algorithms to optimize graph generation from electronic health records for enhanced predictive modeling of multiple chronic conditions. Our main contributions are:
@@ 2 Proposed Methodology
We propose a novel three-part framework that integrates a graph variational autoencoder (GVAE), a Laplacian regularized graph neural network (LR-GNN), and a contextual Bandit (CB) for optimal graph selection. This integration aims to refine and enhance graph representation before training the GNN model, addressing the importance of graph structure in GNN performance [33]. Our approach leverages GVAEs to represent graph features in a latent space, where node representations are functions of connectivity and features (detailed in Section 2.1). GVAE uses the dot product of latent representations to assess node similarity [34, 35], removing edges with low similarity to improve the graph structure [29].
@@
In summary, our methodology combines GVAE and LR-GNN with bandit-inspired optimization to identify and refine graph structures (see Fig.1). This approach aims to improve the quality of graph-based learning and enhance the robustness and accuracy of the GNN model.
Figure 1: Schematic representation of graph variational autoencoder (GVAE) + regularized graph neural network (LR-GNN): The process starts with a fully connected graph defined by its adjacency matrix $A$ and feature set $X$. The GVAE generates various graph variants, which are then used by the LR-GNN to train and evaluate MCC prediction accuracy. Contextual Bandit (CB) is used to identify the best performing graph, and this iterative process is repeated for comprehensive optimization across $m$ iterations.
@@ 2.1 Graph Variational Autoencoder
Our proposed methodology specifically employs a GVAE to learn efficient representations of graph data. By encoding the graph into a lower-dimensional latent space, and by incorporating GVAE, we capture essential information such as features and relationships between nodes (see Fig. 2), thereby reducing the graph’s complexity [36, 37], retaining only the most significant necessary features for subsequent tasks such as node classification using GNNs. The process leverages the power of autoencoders to identify and extract these key features, which can include node similarity, node clustering, and other topological characteristics that are fundamental to understanding the structure of the graph and complex dynamics. In addition, sampling in the latent space allows for the generation of a graph to better represent the EHRs.
@@
Figure 2: Schematic Representation of Graph Variational Autoencoder (GVAE) Architecture. This figure illustrates the GVAE process, where the adjacency matrix A and feature information X are used as inputs. The encoder applies graph convolutional layers to compress the graph information into a latent space representation, represented by mean $\mu$ and variance $\sigma$. Random sampling is performed in the latent space to generate latent variable z. The decoder then reconstructs the graph structure $\hat{A}$. Finally, controlled Gaussian noise is added to the reconstructed graph to generate variants for further processing by the LR-GNN
@@ 2.1.1 Sampling in Latent Space and Kullback-Leibler (KL) Divergence
The encoder of the GVAE computes the mean ($\mu$) and the log-variance ($\log(\sigma^{2})$) for each node in the graph. These parameters define a Gaussian distribution from which the latent representations, $z$ (Eq. 1), are sampled from. This sampling mechanism introduces necessary randomness into the model, reflecting the stochastic nature of real-world graphs. The mathematical representation is as follows:
@@
Here, we use the (batch) stochastic sampling of the latent space to generate diverse and representative latent features that capture the underlying data distribution, improving the model’s ability to generalize across different graph structures [38, 39]. In addition to sampling, KL divergence is utilized as a regularizer in the GVAE latent space. Quantifies the divergence between the learned distribution, characterized by the mean $\mu$, and log-variance, $\log(\sigma^{2})$, of the latent representations, and a prior distribution, typically a standard Gaussian [40]. The KL divergence is computed as follows:
@@
Where $N$ is the dimension of the latent space for the summation to extend the overall dimensions of the latent space. This formulation not only encourages the model to learn meaningful and diverse representations but also ensures that these representations adhere closely to the structure of the prior distribution. By minimizing the KL divergence, the GVAE is guided towards a balance between fitting the data and maintaining a probabilistically sound representation in the latent space. This balance is crucial for the robustness and generalization capabilities of the model, especially in complex graph structures where the underlying data distribution can be intricate and multifaceted [41].
@@ 2.1.2 Edge Adjustment in Decoding
where, $\mathcal{N}(0,\sigma^{2})$ represents Gaussian noise. This technique simulates real-world uncertainties and helps in evaluating the adaptability of the model to structural changes in the graph data. The aim is to ensure that the model remains accurate and stable under different graph conditions, enhancing its utility in practical applications.
@@ 2.3 Graph Neural Network
The overall architecture of the Laplacian Regularized Graph Neural Network (LR-GNN), including the input of different graph variants $\hat{A}$ generated from the GVAE and the process of selecting the best-performing graph, is described in Fig. 3. Each graph variant’s accuracy is determined in terms of predicting MCC. The accuracy of each graph is stored, and once all graphs are analyzed, the data is fed into the $\varepsilon$-G, MAB, and CB models to select the best-performing graph. This selected graph is then used as input for the GVAE to generate a new set of graph variants, repeating the procedure for a fixed number of iterations to enhance the model’s performance.
@@ 2.3.1 Laplacian Regularization
The Laplacian regularization technique is integrated into our GNN model to enhance the learning process [29]. This method involves the construction of a Laplacian matrix $L$, from the graph’s adjacency matrix, $A$, and degree matrix $D$, where $L=D-A$. The regularization term is computed as follows:
@@
Figure 3: Schematic Representation of Laplacian Regularized Graph Neural Network (LR-GNN) Architecture. This figure illustrates the process of generating graph variants, $\hat{A}$ from the GVAE, evaluating their accuracy in predicting MCC and using reinforcement learning models ($\varepsilon$-G, MAB, CB) to select the best-performing graph for iterative refinement.
@@ 3 Numerical Results
In this section, we present the numerical results and analysis of our proposed framework. The primary objective of this section is to evaluate the performance of the GVAE-GNN model using the The Cameron County Hispanic Cohort (CCHC) [49] (Committee for the Protection of Human Subjects: IRB HSC-SPH-03007 Bd) dataset. We evaluate the model’s efficacy and efficiency by conducting a series of experiments to compare the reinforcement learning methods employed, $\varepsilon$-G, MAB, and CB. Through these experiments, we aim to identify the optimal approach for predictive modeling of MCC. The following subsections provide detailed descriptions of the study population, results analysis, and statistical findings.
@@ 3.1 Study population
In this study, we use the CCHC dataset to validate our proposed GNN model. The CCHC is a cohort study that began in 2004, consists primarily of Mexican Americans (98%), and currently enrolls 5,020 patients. It provides a unique opportunity to study health disparities in a community along the Texas-Mexico border. The CCHC employs a rotating recruitment strategy and uses random selection to enroll participants from this community, which has notable health disparities. The inclusion criteria for our model required participants to have valid information on risk variables such as age, body mass index, typical blood pressure, cholesterol level, smoking habits, and family medical history. For model validation, we are focused on a subset of 1,592 patients that met the criteria of not having missing data. This selection aids in analyzing the relationship between five chronic conditions: diabetes, obesity, cognitive impairment, hyperlipidemia, and hypertension.
@@ 3.2 Results Analysis
To validate our model, we conducted three different tests by varying the number of iterations to perform to train the model (i.e., 25, 20, and 15 iterations), to determine the optimal stopping point for evaluating the performance of three reinforcement learning methods: $\varepsilon$-G, MAB, and CB. These tests provided substantial insights into the performance of our integrated GVAE-GNN model for the prediction of MCC.
We initially chose 25 iterations as a starting point to provide a robust baseline for model evaluation. To explore the effects of iteration on model training, we reduced the number of iterations required for training by five, from 25 to 20. This training iteration reduction aimed to identify whether a smaller number of iterations could yield similar performance, potentially reducing computational resources and time. We performed similar operations one more time (15 iterations), further reducing the iteration from 15 didn’t result in competitive results. Thus we will not be reporting this in this work.
The initial set of 25 iterations demonstrated the superior performance of the CB method, as evidenced by its highest average and maximum accuracy, as shown in Fig. 5. In this figure, we also plot the 95% confidence interval for each method based on six iterations to compare the possible worst-best scenarios of each iteration.
Upon reducing the iterations to 20 and then 15 (see Fig.6), we aimed to identify an early stopping point without compromising the performance of the model. The results of the 20 iterations were remarkably close to those of the 25 iterations, suggesting that 20 iterations might be sufficient for a reliable performance evaluation. Similarly to the 25 iterations, the 20 and 15 iterations figures also include the 95% confidence intervals based on six iterations for each method.
@@ 3.3 Statistical Analysis
The statistical analysis of the methods across different iterations provides deeper insights into their performance and reliability.
@@
The $\varepsilon$-G approach demonstrated potential, with an accuracy of 90.06% for 25 iterations, 88.62% in the seventh iteration for 20 and 83.04% for 15 iterations. However, the inherent randomness in the $\varepsilon$-G method, which sometimes favors exploration over exploitation, does not guarantee the best outcome in every iteration. This randomness, while useful in the early stages of a broad exploration, proved to be a limitation when stability and reliability are essential.

### v2

@@ 1 Introduction
The study of multiple chronic conditions (MCC) has extensively explored the impact of various risk factors on MCC networks through advanced statistical and machine learning models. Latent regression Markov clustering has identified transitions between chronic conditions using large datasets from the Department of Veteran Affairs [18]. Temporal Bayesian networks have mapped MCC emergence and patient-level risk factors, utilizing longest-path algorithms to reveal likely comorbidity sequences [19]. Functional continuous-time Bayesian networks have modeled the influence of external variables and lifestyle behaviors on MCC development [20, 21]. Bayesian networks with tree-augmented naive Bayes algorithms have further identified risk factors for hepatocellular carcinoma after hepatectomy [22]. These approaches illustrate the potential of advanced methodologies to elucidate the complex dynamics driving MCC.
@@
Recent advances in the application of machine learning to healthcare have demonstrated significant potential in disease progression prediction and treatment optimization. Carvajal-Rico et al. [34] have further advanced the application of GNNs in healthcare by incorporating Laplacian regularization. This technique enhances the learning process by ensuring similar patient nodes have similar representations, thereby improving the prediction accuracy of models dealing with MCC. Baucum et al. [35] introduced transitional variational autoencoders (tVAEs) to improve reinforcement learning (RL) agents’ performance by generating realistic patient condition trajectories, specifically enhancing the training of medication dosage policies through on-policy RL in intensive care settings. Kmetzsch et al. [36] applied Supervised VAE to predict disease progression scores using multimodal imaging and microRNA data, demonstrating improved accuracy and interpretability in patient outcome predictions. Moreover, Sun et al. [37] highlighted the challenges of using machine learning for disease prediction due to the reliance on abundant manually labeled EMR data, which is often insufficient for rare diseases. Their innovative model leverages external knowledge bases to augment the EMR data and uses GNNs to create highly representative node embeddings for patients, diseases, and symptoms. By aggregating information from connected neighbor nodes, their neural graph encoder can generate embeddings that capture knowledge from both data sources, allowing accurate prediction of general and rare diseases.
@@
In this work, we propose an innovative framework combining a GVAE, a Laplacian regularized graph neural network (LR-GNN), and contextual bandit (CB) algorithms to optimize graph generation from electronic health records for enhanced predictive modeling of multiple chronic conditions. Our main contributions are:
@@ 2 Proposed Methodology
We propose a novel framework that integrates a graph variational autoencoder (GVAE), a LR-GNN, and a CB for optimal graph selection. This integration aims to refine and enhance graph representation before training the GNN model, addressing the importance of graph structure in GNN performance [38]. Our approach leverages GVAEs to represent graph features in a latent space, where node representations are functions of connectivity and features (detailed in Section 2.1). GVAE uses the dot product of latent representations to assess node similarity [39, 40], removing edges with low similarity to improve the graph structure [34].
@@
In summary, our methodology combines GVAE and LR-GNN with bandit-inspired optimization to identify and refine graph structures (see Fig.1(i)). This approach aims to improve the quality of graph-based learning and enhance the robustness and accuracy of the GNN model.
Figure 1: (i) Schematic diagram of the overall proposed model. The process starts with a fully connected graph defined by its adjacency matrix $A$ and feature set $X$. The GVAE generates various graph variants, which are then used by the LR-GNN to train and evaluate MCC prediction accuracy. CB is used to identify the best performing graph, and this iterative process is repeated for comprehensive optimization across $m$ iterations. (ii) The GVAE Architecture, where the adjacency matrix A and feature information X are used as inputs. The encoder applies graph convolutional layers to compress the graph information into a latent space representation, represented by mean $\mu$ and variance $\sigma$. Random sampling is performed in the latent space to generate latent variable z. The decoder then reconstructs the graph structure $\hat{A}$. Finally, controlled Gaussian noise is added to the reconstructed graph to generate variants for further processing by the LR-GNN. (iii) LR-GNN generates graph variants, $\hat{A}$ from the GVAE, evaluating their accuracy in predicting MCC and using reinforcement learning models ($\varepsilon$-G, MAB, CB) to select the best-performing graph for iterative refinement, and finally (iv) reinforcement learning framework.
@@ 2.1 Graph Variational Autoencoder
Our proposed methodology specifically employs a GVAE to learn efficient representations of graph data. By encoding the graph into a lower-dimensional latent space, and by incorporating GVAE, we capture essential information such as features and relationships between nodes (see Fig.1(ii)), thereby reducing the graph’s complexity [41, 42], retaining only the most significant necessary features for subsequent tasks such as node classification using GNNs. The process leverages the power of autoencoders to identify and extract these key features, which can include node similarity, node clustering, and other topological characteristics that are fundamental to understanding the structure of the graph and complex dynamics. In addition, sampling in the latent space allows for the generation of a graph to better represent the EHRs.
@@ 2.1.1 Sampling in Latent Space and Kullback-Leibler (KL) Divergence
The encoder of the GVAE computes the mean ($\mu$) and the log-variance ($\log(\sigma^{2})$) for each node in the graph. These parameters define a Gaussian distribution from which the latent representations, $z$ (Eq. 1), are sampled. This sampling mechanism introduces necessary randomness into the model, reflecting the stochastic nature of real-world graphs. The mathematical representation is as follows:
@@
Batch stochastic sampling of the latent space to generate diverse and representative latent features that capture the underlying data distribution, improving the model’s ability to generalize across different graph structures [43, 44]. In addition to sampling, KL divergence is utilized as a regularizer in the GVAE latent space. KL divergence, often described as ”a measure of how one probability distribution differs from another,” ensures that the latent representations generated by the GVAE remain probabilistically meaningful. In this case, it quantifies the divergence between the learned distribution, characterized by the mean $\mu$, and log-variance, $\log(\sigma^{2})$, of the latent representations, and a prior distribution, typically a standard Gaussian [45]. The KL divergence is computed as follows:
@@
where $N$ is the dimension of the latent space for the summation to extend the overall dimensions of the latent space. This formulation not only encourages the model to learn meaningful and diverse representations but also ensures that these representations adhere closely to the structure of the prior distribution. By minimizing the KL divergence, the GVAE achieves a balance between fitting the data and maintaining a probabilistically sound representation in the latent space. This balance is crucial for the robustness and generalization capabilities of the model, especially when dealing with complex graph structures where the underlying data distribution can be intricate and multifaceted [46].
@@ 2.1.2 Edge Adjustment in Decoding
where, $\mathcal{N}(0,\sigma^{2})$ represents Gaussian noise. This approach introduces controlled variability to simulate real-world uncertainties, enabling the evaluation of the model’s robustness and adaptability to structural changes in graph data. The goal is to ensure that the model maintains both accuracy and stability across diverse graph conditions, thereby increasing its reliability and practicality in real-world scenarios.
@@ 2.3 Graph Neural Network
The overall architecture of the Laplacian Regularized Graph Neural Network (LR-GNN), including the input of different graph variants $\hat{A}$ generated from the GVAE and the process of selecting the best-performing graph, is described in Fig.1(iii). Each graph variant’s accuracy is determined in terms of predicting MCC. The accuracy of each graph is stored, and once all graphs are analyzed, the data is fed into the $\varepsilon$-G, MAB, and CB models to select the best-performing graph. This selected graph is then used as input for the GVAE to generate a new set of graph variants, repeating the procedure for a fixed number of iterations to enhance the model’s performance.
@@ 2.3.1 Laplacian Regularization
The Laplacian regularization technique is integrated into our GNN model to enhance the learning process and improve node classification accuracy [34]. This method leverages the Laplacian matrix $L$, constructed from the graph’s adjacency matrix $A$ and degree matrix $D$, where $L=D-A$, to enforce smoothness in the learned node embeddings. Specifically, nodes with similar structural or feature properties are encouraged to have similar embeddings, which aligns with the inherent relationships in the graph data. The regularization term is defined as:
@@
where $\lambda$ is a regularization parameter, $y_{\text{pred}}$ represents the predicted node embeddings, and $Ly_{\text{pred}}$ captures the influence of neighboring nodes. This term ensures that the embeddings respect the graph’s local structure, leading to more robust predictions. The detailed development and validation of the LR-GNN model, including the design of the Laplacian regularization term, are provided by [34]. In this study by the incorporation of the Laplacian regularization, the GNN model achieves smoother and more reliable embeddings, which contribute significantly to the overall framework’s robustness and accuracy, specifically for node classification.
@@ 3 Numerical Results
In this section, we present the numerical results and analysis of our proposed framework. The primary objective of this section is to evaluate the performance of the GVAE-GNN model using the Cameron County Hispanic Cohort (CCHC) [53] (Committee for the Protection of Human Subjects: IRB HSC-SPH-03007 Bd) dataset. We evaluate the model’s efficacy and efficiency by conducting a series of experiments to compare the reinforcement learning methods employed: $\varepsilon$-G, MAB, and CB. Through these experiments, we aim to identify the optimal approach for predictive modeling of MCC. The CB method was the most effective approach during this process, as it accounts for the contextual information of the graph variants while maintaining a balance between exploration and exploitation. In our implementation, context refers to the feature vectors extracted from each graph variant, which encapsulates structural properties and node attributes for predictive modeling. These features include, node embeddings, and statistical measures that characterize each graph representation. By integrating these contextual features, the CB model dynamically adapts its selection strategy, prioritizing graph variants that are more likely to improve the accuracy of the model. This ability to leverage graph-specific context allows the CB model to make more informed decisions compared to non-contextual approaches, leading to superior predictive performance. The integration of GVAE, RL, and GNN ensures a cohesive flow: (i) GVAE generates graph variants, (ii) GNN trained on the variants generates evaluation metrics for the graphs, and finally, (iii) RL evaluates and selects the best graph variant and the associated robust features for the best performing model. This iterative optimization ensures that the GNN leverages graph structures that maximize its predictive performance while efficiently exploring the graph space. By connecting the GVAE-generated graphs to the bandit optimization process, our framework achieves a seamless integration of generative modeling, optimization, and graph-based learning.
The following subsections provide detailed descriptions of the study population, analysis, and statistical findings.
@@ 3.1 Study population
In this study, we use the CCHC dataset to validate our proposed GNN model. The CCHC is a cohort study that began in 2004, consists primarily of Mexican Americans (98%), and currently enrolls 5,020 patients. It provides a unique opportunity to study health disparities in a community along the Texas-Mexico border. The CCHC employs a rotating recruitment strategy and uses random selection to enroll participants from this community, which has notable health disparities. The inclusion criteria for our model required participants to have valid information on risk variables such as age, body mass index (BMI), typical blood pressure, cholesterol level, smoking habits, parental history of diabetes (father/mother), parental history of hypertension (father/mother), systolic and diastolic blood pressure, mini-mental state examination (MMSE) score (a cognitive impairment indicator), and gender.
For model validation, we are focused on a subset of 1,592 patients that met the criteria of not having missing data. This selection aids in analyzing the relationship between five chronic conditions: diabetes, obesity, cognitive impairment, hyperlipidemia, and hypertension. Each patient (node) in our framework is assigned a feature vector consisting of these risk factors, enabling the model to learn associations between patients’ risk factors and MCC outcomes. The LR-GNN model predicts the possible outcome for a patient by categorizing it into one of 32 possible classes, representing different combinations of the five chronic conditions. This allows the model to capture patterns of MCC rather than predicting single conditions. By leveraging graph-based learning models, we can analyze the complex interaction among risk factors and improve predictive accuracy in identifying high-risk patients.
@@ 3.2 Results Analysis
To validate our model, we conducted three different tests by varying the number of iterations to train the model (i.e., 25, 20, and 15 iterations) to determine the optimal stopping point for evaluating the performance of three reinforcement learning methods: $\varepsilon$-G, MAB, and CB. These tests provided substantial insights into the performance of our integrated GVAE-GNN model for the prediction of MCC.
We initially chose 25 iterations as a starting point to provide a robust baseline for model evaluation. To explore the effects of iteration on model training, we reduced the number of iterations required for training by five, from 25 to 20. This training iteration reduction aimed to identify whether a smaller number of iterations could yield similar performance, potentially reducing computational resources and time. One additional reduction to 15 iterations was performed; however, reducing iterations further from 15 did not yield competitive results. Thus, these results are not reported in this work.
The initial set of 25 iterations demonstrated the superior performance of the CB method, as evidenced by its highest average and maximum accuracy, as shown in Fig. 3. In this figure, we also plot the 95% confidence interval for each method based on six iterations to compare the possible worst-best scenarios for each method.
Upon reducing the iterations to 20 and 15 (see Fig.4), we aimed to identify an early stopping point without compromising the performance of the model. The results of the 20 iterations were remarkably close to those of the 25 iterations, suggesting that 20 iterations might suffice for reliable performance evaluation. Similarly to the 25 iterations, the 20- and 15-iteration figures also include the 95% confidence intervals based on six iterations per method.
@@
3.2.1 Comparison with Classic Machine Learning Models
To further evaluate the efficacy of our GVAE-GNN model, we compared it against three classic machine learning models: K-Nearest Neighbors (KNN), Logistic Regression (LR), and a Feedforward Neural Network (FNN). Consistent preprocessing steps were applied across all methods to ensure fairness in the comparison. The results are summarized below in Table 2:
Table 2: Comparison between best performing set-up and classic non-graph based machine leanring models
Model/Metric GVAE-GNN KNN LR FNN
Test Accuracy 0.8316 0.2720 0.3013 0.2579
F1-Score 0.8238 0.2360 0.2337 0.2202
Recall 0.8240 0.2720 0.3013 0.2579
Precision 0.8288 0.2539 0.2106 0.2022
In contrast, our GVAE-GNN model consistently outperformed all baseline models. Its ability to leverage both structural and feature information contributed to its superior results. By integrating graph-based representations and reinforcement learning, the GVAE-GNN demonstrated greater accuracy, robustness, and stability across varying iterations. As detailed in the comparative analysis, KNN, Logistic Regression, and FNN all faced significant challenges in handling the dataset’s multi-class nature and high feature dimensionality. While these methods offer simplicity and computational efficiency, their inability to capture nuanced relationships in the data led to suboptimal performance. In contrast, the GVAE-GNN combined with the CB approach not only outperformed these models but also demonstrated greater flexibility and adaptability, reinforcing its suitability for MCC prediction.
@@ 3.3 Statistical Analysis
The statistical analysis of the methods across different iterations provides deeper insights into their performance and reliability. The $\varepsilon$-G approach demonstrated potential, with an accuracy of 90.06% for 25 iterations, 88.62% in the seventh iteration for 20 and 83.04% for 15 iterations. However, the inherent randomness in the $\varepsilon$-G method, which sometimes favors exploration over exploitation, does not guarantee the best outcome in every iteration. This randomness, while useful in the early stages of a broad exploration, proved to be a limitation when stability and reliability are essential.
@@
F1-Score 0.7358 0.6615 0.7622 0.7852 0.6758 0.8116 0.7901 0.6633 0.8238
Recall 0.7412 0.6686 0.7636 0.7870 0.6814 0.8123 0.7916 0.6698 0.8240
Precision 0.7549 0.6893 0.7741 0.7960 0.6978 0.8180 0.7968 0.6887 0.8288
AUC 0.9564 0.9403 0.9675 0.9673 0.9392 0.9766 0.9686 0.9364 0.9779